# QSO MUSEUM II: Search for extended Ly emission around 8 quasar pairs

### v1

@@ 5.1 Comparison to previous quasar pair studies
To our knowledge, the number of previous works targeting extended emission on CGM/IGM scales around physically associated quasar pairs is limited to seven studies (Cai et al. 2018; Arrigoni Battaia et al. 2019b; Lusso et al. 2019; Li et al. 2023), three of which have been serendipitous detections since the faint companions were not known during observation planning (Cantalupo et al. 2014; Hennawi et al. 2015; Arrigoni Battaia et al. 2018). These studies unanimously reported Ly nebulae with exceptional extents (up to 460 kpc). In our work however, extended emission often does not cover similarly large areas (¿104 kpc2) or reaches such high luminosities (¿ erg s-1).
@@
Li et al. (2023) targeted quasar pairs at z2.5 with one bright (¡18 mag) quasar and included a known ELAN first published in Cai et al. (2018), resulting in two out of 5 quasar pairs meeting their criteria for ELANe. Compared to that work, nebulae studied here have much smaller extents, even when the integrated luminosity is comparable. This fact is indicative of the differing methods to determine what significant emission means. Here we decided to employ a more conservative approach in order to match previous narrow-band study designs, leading to smaller 2 isophotes. A comparison of these results is therefore not straightforward, but we can firmly state that our data do not unveil nebulosities as bright as those in prototypical ELANe, which have SB erg s-1 cm-2 arcsec-2 on hundreds of kiloparsecs (e.g., Hennawi et al. 2015; Arrigoni Battaia et al. 2018). Therefore, the sample presented here does not confirm the overabundance of ELANe around the overall population of quasar pairs as we do not identify any ELAN within our sample of 8 pairs. Importantly, the targeted sample extends the study of large-scale emission around quasar pairs to additional systems in which both quasars are faint (Fig. 2). This fact is also one of the reasons why we do not detect bright and extended nebulae around all pairs (see more details in Sect. 5.2). In Fig. 11 we compare our stacked quasar pair Ly profile to the one obtained using the same method as employed in this work by Li et al. (2023) for quasar pairs after correcting them for cosmological dimming. This exercise shows that overall the SB levels are consistent, contrary to the results found for Ly emission around single bright quasars, with quasars having a factor of brighter extended Ly emission (Cai et al. 2019). In other words, naively, we would have expected quasar pairs at to be associated with brighter emission than quasar pairs for similar AGN magnitude ranges. Therefore we infer that the emission in our sample is modest because of the average lower luminosity of the targeted quasars (Sect. 5.2).
@@
Contrarily, the Ly nebula published by Lusso et al. (2019) in the MUSE Ultra Deep Field (MUDF) around a quasar pair ( and ) shows a similar morphology to some objects studied here, in particular to quasar 4.1. Both pairs have a projected separation of about 500 physical kpc, the Ly emission around the brighter quasar extending toward the fainter second quasar, and showing an additional smaller nebula along such direction (Fig. 6). While the extremely long integration time of 40 hours in Lusso et al. (2019) allows for sensitivity to exceptionally faint emission, the nebula sizes still do not exceed the expected virial radius. The elongation between that quasar pair also appears to fit the relation shown in Fig. 13 for our systems. Quasar pairs like the one targeted in Lusso et al. (2019) and quasar pair 4 are therefore excellent systems for ultra-deep integrations to unveil intergalactic bridges and directly study the IGM.
@@
Lastly, one object studied in this work, pair 8, has already been published in Arrigoni Battaia et al. (2019b). While most of the results are the same, extended emission is detected in that work using a 3D mask, making it more sensitive to faint and narrow emission. Therefore, the area detected above 2 is bigger and the nebula is more coherent than in this work. Since the velocity information in that work is computed using only the wavelength ranges within the 3D mask and included the nebula outskirts, the average velocity dispersion resulted to be much lower (162 km s-1). We caution that the dispersion calculation is highly dependant on the spectral width considered, and the inclusion of spaxels spanning 1.25 to 2.5 in the 3D mask will necessarily lead to very low dispersion measures. This comparison highlights the importance of considering the different methods of extended emission extraction when studying Ly nebula kinematics.
@@ 5.2 Comparison to single quasar samples
A large number of single quasar fields around redshift 3 have already been targeted with IFUs to study extended Ly emission, although mostly focusing on more luminous quasars than presented in this work (Fig. 2). Multiple surveys comprising more than one hundred bright quasars detected extended Ly emission on scales of, on average, 80 to 150 kpc in maximum projected size around 100 % of the targets (Borisova et al., 2016; Arrigoni Battaia et al., 2019a; Fossati et al., 2021). This is in stark contrast to the quasar pair sample presented here, where two of the 14 observed quasars are not associated with extended emission and the detected nebulae are comparatively small. This difference can be appreciated by comparing the stacked SB profiles presented in Fig. 11. The SB profile of our quasar pair sample is 5 times fainter than the SB profile around bright single quasars. This difference is remarkable as some of the targeted quasar pairs could trace ongoing mergers between massive galaxies. In such events several processes (e.g., tides) could boost the gas densities on CGM scales and hence the associated Ly emission (Sect. 5.4). Further, the differences cannot be explained by the observational setup as the same method is adopted for most of the targets in the aforementioned studies (i.e., an exposure time of at least 1 hour in MUSE WFM) but it could possibly arise due to target selection as our quasar pair sample has not been selected from SDSS based on having bright quasars in the pair.
Indeed, a single quasar survey selecting 12 similarly low luminosity quasars at found extended emission in each targeted field with a minimum extent of 60 pkpc, but did show a connection between quasar Ly peak luminosity and nebula luminosity (Mackenzie et al., 2021). We note that Mackenzie et al. (2021) adopted an optimal extraction method for constructing a 3D region for the study of the Ly emission. This technique, applied channel by channel, is more sensitive to the fainter portions of the nebulae with respect to our method. For this reason we tested our analysis by applying it to the two extreme nebula cases in that sample (see Appendix B). We find that our analysis is able to reproduce the principal shape and luminosity of the brightest nebula reported in that paper, while our measurement of the extent deviates by 30 %. The faint nebula is only detected in the peak of the nebula emission and thus is 3 times smaller in maximum extent than reported in Mackenzie et al. (2021). Therefore, the difference in maximum extent of 15 % between the faint single-quasar sample and our sample of quasar pairs is likely in part due to the difference in methodology.
@@
Further evidence of the presence of large-scale structures comes from the morphologies of the discovered nebulae around quasar pairs. In section 4.3 we have shown a comparison of their morphologies (described by their axis ratios) with respect to nebulae around single quasars (Fig. 12). Morphologies of nebulae around single quasars are mostly circular (Borisova et al., 2016; Arrigoni Battaia et al., 2019a), with even higher symmetry for faint quasars (Mackenzie et al., 2021). ELANe, known to host multiple AGNs, are more asymmetric (Arrigoni Battaia et al. 2023a). Similarly, quasar pair nebulae have a median axis ratio lower than nebulae around single quasars (Fig. 12). This is akin to redshift 2 single quasar nebulae with a median value of (Cai et al., 2019), hypothesized to originate in galaxy mergers, which peak at . While this interpretation can be supported in the context of quasar pairs, the spread of in our small sample is substantial and is not directly correlated with the projected distance. Asymmetries can therefore not only be driven by galaxy mergers, but also by intergalactic structures.
@@ 5.3 Direction of nebulae with respect to the quasar pairs
This expectation is based on the assumption that a nebula semimajor axis (used to compute ) is a good proxy for the direction of the underlying cool gas distribution. Cosmological simulations post-processed with a radiative transfer tool have shown that this is true if the quasar host galaxy is not edge-on (Costa et al., 2022). The more edge-on the host galaxy the more lopsided the extended Ly emission. The fact that the majority of quasar nebulae are symmetric may therefore indicates that their host galaxies are far from being edge-on and are well sampling the surrounding hydrogen gas distribution (Arrigoni Battaia et al., 2023a).
These simple predictions are further complicated by the fact that the morphology of the Ly nebulae (and hence and the angle ) can be influenced by a multitude of factors. Indeed it has been shown that the following factors all play a role in shaping the extents and luminosities of quasar nebulae: dust in the host galaxy or the CGM (Smith et al., 2022; González Lobos et al., 2023) as dust can efficiently absorb Ly photons; the molecular gas mass of the host galaxy, being a proxy for denser and dust-richer galaxies which may prevent the escape of ionizing and Ly photons (Muñoz-Elgueta et al., 2022); the Ly luminosity and/or magnitude of the powering sources, with the trend of brighter quasars having brighter and often more extended nebulae (Mackenzie et al., 2021); the availability of cool, 104 K gas that has the potential to be Ly bright; and geometry effects (e.g., orientation of the quasar’s ionizing cones with respect to the gas distribution; Obreja et al. 2024; orientation of cosmic web filaments with respect to the observing angle, which could lead to absorption of Ly). In addition to those, quasar pair nebulae might not trace exclusively inflowing CGM gas (see Sect. 5.4), especially in close pairs.
@@ 5.4 Origin of the cool gas probed by Ly around quasar pairs
The gas traced by extended Ly emission around quasars is usually assumed to be the cool ( K) phase of the CGM, supported by typical Ly halo radii kpc for the bulk of the emission, just short of the expected virial radius for quasar halos (e.g., Arrigoni Battaia et al. 2019a). This is further confirmed by the levels of the Ly SB profiles around quasars as a function of redshift. The absence of an evolution from down to (Farina et al. 2019; Fossati et al. 2021) and the subsequent decrease to (Arrigoni Battaia et al. 2019a; Cai et al. 2019) seems in agreement with the theoretical expectation (e.g., Dekel & Birnboim 2006) that high- () massive halos are able to accrete gas in a cool phase keeping the Ly emission roughly constant, while at lower redshift the gas mass in the cool phase, and hence the Ly emission, is reduced by shock heating. At moderate depths ( erg s-1 cm-2 arcsec-2), few rare exceptions show very extended emission that can exceed the projected virial radius of the host halo and therefore reach into the IGM (e.g., Cantalupo et al. 2014; Arrigoni Battaia et al. 2019a). However, most of these outliers have active companions.
@@
Quasar pairs seem to follow the same trends as single quasars, but with, on average, enhanced Ly emission at larger distances (Sect. 5.2). The Ly emission in quasar pairs can therefore better trace IGM structures given the closer presence of an additional massive halo. Indeed, cosmological simulations report the presence of denser and thicker filaments in denser environments (Cautun et al. 2014), similar to the locations quasar pairs are expected to populate in the cosmic web (e.g., Boris et al. 2007; Onoue et al. 2018). Given the presence of the second quasar, they are at least in a group environment. For this reason, processes favored by dense environments could affect the cool gas distribution more than around single quasars. In particular, galaxy interactions may become especially important in pairs at close separation. Both tidal stripping and ram pressure stripping could significantly enhance the density of the CGM in a preferential direction aligned with an interacting pair (Salem et al., 2015; Samuel et al., 2023) and therefore may boost the observed surface brightness out to large radii. Evidence for gas stripping around high- quasars is indeed mounting (e.g., Chen et al. 2021; Vayner et al. 2023). The effects of these environmental processes can, in part, explain the alignment of extended nebulae, specifically in close pairs (Fig. 13), and the comparatively flat surface brightness profile observed in the median stack of quasar pairs (Fig. 11) and especially apparent in the profiles of some individual quasar pairs, for example pair 6 (Fig. 8). However, the properties of the extended emission studied here (Table 4) do not correlate with the angular separation between the quasar pairs, as would be expected if the galaxy interaction alters the CGM in an observable way. Therefore, the morphology of extended Ly emission alone does not provide evidence for the aforementioned environmental effects. Summarizing, the diversity of extended Ly emission around quasar pairs probes the combined effects of dense environments and close massive halos, possibly tracing the surrounding large scale structure.
@@ 6 Summary and conclusions
Motivated by previously discovered ELANe around quasar pairs and signs of filamentary gas connecting the two quasars in multiple systems, we have conducted the largest study of extended Ly emission around quasar pairs to date. Our sample consists of 14 quasars in 8 pairs observed with VLT/MUSE with a snapshot strategy (45 minutes/source). The targets span redshifts and -band magnitudes between 18 and 22.8, dimmer than most single-quasar samples. Our key findings are as follows:
@@
12 out of the 14 observed quasars are associated with extended Ly emission spanning 40 to 160 kpc, and 3 of 5 close pairs are embedded in the same Ly nebula (Table 4, Figs. 3 - 10). In contrast to previously discovered quasar pairs, none of these nebulosities meet the criteria for ELANe in either or extent, lowering the fraction of ELANe around known quasar pairs significantly (Sect. 5.1). On average, the nebulae in our sample are 15 % smaller in maximum projected extent than nebulae around similar single-quasar samples, but have comparably quiescent kinematics (Section 5.2).
@@
Larger nebulae are preferentially aligned with the expected direction of the large-scale structure between the associated quasar pair, while smaller nebulae tend to be misaligned (Fig. 13). This fact is indicative of a direct connection between filaments and the cool CGM (Sect. 5.3). Wide pairs highly aligned with their filaments, like quasar pair 4, are excellent targets for ultra-deep follow-up observations to detect these faint IGM structures.
@@
The SB profile of quasar pairs is comparable in brightness to single-quasar profiles at similar redshift, but is fainter close to the quasar and falls off more gradually with a power law slope of (Fig. 11). Due to galaxy interactions important in close quasar pairs, the higher surface brightness at larger radii can in part be explained by stripped gas increasing the densities in the CGM. The presence of a second large halo can also add to the Ly emission through increased contribution from IGM structures (Sect. 5.4).
@@
Extended emission around this sample of quasar pairs is morphologically diverse, but on average more asymmetric than single-quasar nebulae with large offsets between quasar position and flux-weighted nebula centroid (Table 4, Fig. 12). This can be explained by the presence of more sub-structures in the overdense regions likely populated by quasar pairs and by close pairs sometimes sharing the same nebula.
@@
We do not find a trend in brightness or morphology of the extended Ly emission with the projected distance between the associated quasar pair, as would be expected if the host galaxies’ interaction influences the CGM in a way observable through the extended Ly emission (Sect. 5.4). While the sample studied in this work only encompasses close ( pkpc) and wide ( pkpc) pairs, extending the sample of observed quasar pairs to intermediate projected distances has the potential to reveal the full picture.
@@
This study highlights once again the promise of Ly nebulae around quasars tracing the large-scale structure in dense environments. As quasar pair nebulae show a higher surface brightness level at larger radii than individual quasars, they provide the most accessible laboratory to directly study the IGM and its connection to the associated galaxies. Indeed, our sample contains excellent targets for future deep and ultra-deep observations to study the large-scale gaseous structures of the IGM (e.g., Tornotti et al. 2024). Additional snapshot observations of quasar pairs are needed to increase the number of such suitable targets and therefore cover the diversity of IGM filament structures and active galaxy interactions at different separation stages.

### v2

@@ 5.1 Comparison to previous quasar pair studies
To our knowledge, the number of previous works targeting extended emission on CGM/IGM scales around physically associated quasar pairs is limited to seven studies (Cai et al. 2018; Arrigoni Battaia et al. 2019b; Lusso et al. 2019; Li et al. 2023), three of which have been serendipitous detections since the faint companions were not known during observation planning (Cantalupo et al. 2014; Hennawi et al. 2015; Arrigoni Battaia et al. 2018). These studies unanimously reported Ly nebulae with exceptional extents (up to 460 kpc). In our work, however, extended emission often does not cover similarly large areas (¿104 kpc2) or reach such high luminosities (¿ erg s-1).
@@
Li et al. (2023) targeted quasar pairs at with one bright (¡18 mag) quasar and included a known ELAN first published in Cai et al. (2018), resulting in two out of five quasar pairs meeting their criteria for ELANe. Compared to that work, nebulae studied here have much smaller extents, even when the integrated luminosity is comparable. This fact is indicative of the differing methods to determine what significant emission means. Here we decided to employ a more conservative approach in order to match previous narrowband study designs, leading to smaller 2 isophotes. A comparison of these results is therefore not straightforward, but we can firmly state that our data do not unveil nebulosities as bright as those in prototypical ELANe, which have SB erg s-1 cm-2 arcsec-2 on hundreds of kiloparsecs (e.g., Hennawi et al. 2015; Arrigoni Battaia et al. 2018). Therefore, the sample presented here does not confirm the overabundance of ELANe around the overall population of quasar pairs as we do not identify any ELAN within our sample of 8 pairs. Importantly, the targeted sample extends the study of large-scale emission around quasar pairs to additional systems in which both quasars are faint (Fig. 2). This fact is also one of the reasons why we do not detect bright and extended nebulae around all pairs (see more details in Sect. 5.2). In Fig. 4 we compare our stacked quasar pair Ly profile to the one obtained using the same method as employed in this work by Li et al. (2023) for quasar pairs after correcting them for cosmological dimming. This exercise shows that overall the SB levels are consistent, contrary to the results found for Ly emission around single bright quasars, with quasars having a factor of brighter extended Ly emission (Cai et al. 2019). In other words, naively, we would have expected quasar pairs at to be associated with brighter emission than quasar pairs for similar AGN magnitude ranges. Therefore we inferred that the emission in our sample is modest because of the average lower luminosity of the targeted quasars (Sect. 5.2).
@@
Contrarily, the Ly nebula published by Lusso et al. (2019) in the MUSE Ultra Deep Field (MUDF) around a quasar pair ( and ) shows a similar morphology to some objects studied here, in particular to quasar 4.1. Both pairs have a projected separation of about 500 physical kiloparsecs, with the Ly emission around the brighter quasar extending toward the fainter second quasar and showing an additional smaller nebula along the same direction (Fig. 9). While the extremely long integration time of 40 hours in Lusso et al. (2019) allows for sensitivity to exceptionally faint emission, the nebula sizes still do not exceed the expected virial radius. The elongation between that quasar pair also appears to fit the relation shown in Fig. 6 for our systems. Quasar pairs like the one targeted in Lusso et al. (2019) and quasar pair 4 are therefore excellent systems for ultra-deep integrations to unveil intergalactic bridges and directly study the IGM.
@@
Lastly, one object studied in this work, pair 8, has already been published in Arrigoni Battaia et al. (2019b). While most of the results are the same, extended emission is detected in that work using a 3D mask, making it more sensitive to faint and narrow emission. Therefore, the area detected above 2 is bigger, and the nebula is more coherent than in this work. Since the velocity information in that work was computed using only the wavelength ranges within the 3D mask and included the nebula outskirts, the average velocity dispersion is much lower (162 km s-1). We caution that the dispersion calculation is highly dependent on the spectral width considered, and the inclusion of spaxels spanning 1.25 to 2.5 in the 3D mask necessarily leads to very low dispersion measures. This comparison highlights the importance of considering the different methods of extended emission extraction when studying Ly nebula kinematics.
@@ 5.2 Comparison to single quasar samples
A large number of single quasar fields around redshift three have already been targeted with IFUs to study extended Ly emission, although mostly focusing on more luminous quasars than presented in this work (Fig. 2). Multiple surveys comprising more than one hundred bright quasars detected extended Ly emission on scales of, on average, 80 to 150 kpc in maximum projected size around 100 % of the targets (Borisova et al., 2016; Arrigoni Battaia et al., 2019a; Fossati et al., 2021). This is in stark contrast to the quasar pair sample presented here, where two of the 14 observed quasars are not associated with extended emission and the detected nebulae are comparatively small. This difference can be appreciated by comparing the stacked SB profiles presented in Fig. 4. The SB profile of our quasar pair sample is 5 times fainter than the SB profile around bright single quasars. This difference is remarkable as some of the targeted quasar pairs could trace ongoing mergers between massive galaxies. In such events several processes (e.g., tides) could boost the gas densities on CGM scales and hence the associated Ly emission (Sect. 5.4). Further, the differences cannot be explained by the observational setup as the same method is adopted for most of the targets in the aforementioned studies (i.e., an exposure time of at least 1 hour in MUSE WFM) but it could possibly arise due to target selection as our quasar pair sample has not been selected from SDSS based on having bright quasars in the pair.
Indeed, a single quasar survey selecting 12 similarly low luminosity quasars at found extended emission in each targeted field with a minimum extent of 60 pkpc, but did show a connection between quasar Ly peak luminosity and nebula luminosity (Mackenzie et al., 2021). We note that Mackenzie et al. (2021) adopted an optimal extraction method for constructing a 3D region for the study of the Ly emission. This technique, applied channel by channel, is more sensitive to the fainter portions of the nebulae with respect to our method. For this reason we tested our analysis by applying it to the two extreme nebula cases in that sample (see Appendix C). We find that our analysis is able to reproduce the principal shape and luminosity of the brightest nebula reported in that paper, while our measurement of the extent deviates by 30 %. The faint nebula is only detected in the peak of the nebula emission and is thus three times smaller in maximum extent than reported in Mackenzie et al. (2021). Therefore, the difference in maximum extent of 15 % between the faint single-quasar sample and our sample of quasar pairs is likely in part due to the difference in methodology.
@@
Further evidence of the presence of large-scale structures comes from the morphologies of the discovered nebulae around quasar pairs. In section 4.3 we have shown a comparison of their morphologies (described by their axis ratios) with respect to nebulae around single quasars (Fig. 5). Morphologies of nebulae around single quasars are mostly circular (Borisova et al., 2016; Arrigoni Battaia et al., 2019a), with even higher symmetry for faint quasars (Mackenzie et al., 2021). ELANe, known to host multiple AGNs, are more asymmetric (Arrigoni Battaia et al. 2023a). Similarly, quasar pair nebulae have a median axis ratio lower than nebulae around single quasars (Fig. 5). This is akin to redshift 2 single quasar nebulae with a median value of (Cai et al., 2019) hypothesized to originate in galaxy mergers, which peak at . While this interpretation can be supported in the context of quasar pairs, the spread of in our small sample is substantial and is not directly correlated with the projected distance. Asymmetries can therefore not only be driven by galaxy mergers, but also by intergalactic structures.
@@ 5.3 Direction of nebulae with respect to the quasar pairs
This expectation is based on the assumption that a nebula semi-major axis (used to compute ) is a good proxy for the direction of the underlying cool gas distribution. Cosmological simulations post-processed with a radiative transfer tool have shown that this is true if the quasar host galaxy is not edge-on (Costa et al., 2022). The more edge-on the host galaxy the more lopsided the extended Ly emission. The fact that the majority of quasar nebulae are symmetric may therefore indicate that their host galaxies are far from being edge-on and are sampling the surrounding hydrogen gas distribution well (Arrigoni Battaia et al., 2023a).
These simple predictions are further complicated by the fact that the morphology of the Ly nebulae (and hence and the angle ) can be influenced by a multitude of factors. Indeed it has been shown that the following factors all play a role in shaping the extents and luminosities of quasar nebulae: dust in the host galaxy or the CGM (Smith et al., 2022; González Lobos et al., 2023) as dust can efficiently absorb Ly photons; the molecular gas mass of the host galaxy, being a proxy for denser and more dust rich galaxies that may prevent the escape of ionizing and Ly photons (Muñoz-Elgueta et al., 2022); the Ly luminosity and/or magnitude of the powering sources, with the trend of brighter quasars having brighter and often more extended nebulae (Mackenzie et al., 2021); the availability of cool, 104 K gas that has the potential to be Ly bright; and geometry effects (e.g., orientation of the quasar’s ionizing cones with respect to the gas distribution; Obreja et al. 2024; orientation of cosmic web filaments with respect to the observing angle, which could lead to absorption of Ly). In addition to those, quasar pair nebulae might not trace exclusively inflowing CGM gas (see Sect. 5.4), especially in close pairs.
@@ 5.4 Origin of the cool gas probed by Ly around quasar pairs
The gas traced by extended Ly emission around quasars is usually assumed to be the cool ( K) phase of the CGM, supported by typical Ly halo radii kpc for the bulk of the emission, just short of the expected virial radius for quasar halos (e.g., Arrigoni Battaia et al. 2019a). This is further confirmed by the levels of the Ly SB profiles around quasars as a function of redshift. The absence of an evolution from down to (Farina et al. 2019; Fossati et al. 2021) and the subsequent decrease to (Arrigoni Battaia et al. 2019a; Cai et al. 2019) seems in agreement with the theoretical expectation (e.g., Dekel & Birnboim 2006) that high- () massive halos are able to accrete gas in a cool phase keeping the Ly emission roughly constant, while at lower redshift, the gas mass in the cool phase (and hence the Ly emission) is reduced by shock heating. At moderate depths ( erg s-1 cm-2 arcsec-2), a few rare exceptions show very extended emission that can exceed the projected virial radius of the host halo and therefore reach into the IGM (e.g., Cantalupo et al. 2014; Arrigoni Battaia et al. 2019a). However, most of these outliers have active companions.
@@
Quasar pairs seem to follow the same trends as single quasars, but with, on average, enhanced Ly emission at larger distances (Sect. 5.2). The Ly emission in quasar pairs can therefore better trace IGM structures given the closer presence of an additional massive halo. Indeed, cosmological simulations report the presence of denser and thicker filaments in denser environments (Cautun et al. 2014), similar to the locations quasar pairs are expected to populate in the cosmic web (e.g., Boris et al. 2007; Onoue et al. 2018). Given the presence of the second quasar, they are at least in a group environment. For this reason, processes favored by dense environments could affect the cool gas distribution more than around single quasars. In particular, galaxy interactions may become especially important in pairs at close separation. Both tidal stripping and ram pressure stripping could significantly enhance the density of the CGM in a preferential direction aligned with an interacting pair (Salem et al., 2015; Samuel et al., 2023) and may therefore boost the observed surface brightness out to large radii. Evidence for gas stripping around high- quasars is indeed mounting (e.g., Chen et al. 2021; Vayner et al. 2023). The effects of these environmental processes can, in part, explain the alignment of extended nebulae, specifically in close pairs (Fig. 6), and the comparatively flat surface brightness profile observed in the median stack of quasar pairs (Fig. 4) and especially apparent in the profiles of some individual quasar pairs, for example pair 6 (Fig. 11). However, the properties of the extended emission studied here (Table 3) do not correlate with the angular separation between the quasar pairs, as would be expected if the galaxy interaction alters the CGM in an observable way. Therefore, the morphology of extended Ly emission alone does not provide evidence for the aforementioned environmental effects. Summarizing, the diversity of extended Ly emission around quasar pairs probes the combined effects of dense environments and close massive halos, possibly tracing the surrounding large scale structure.
@@ 6 Summary and conclusions
Motivated by previously discovered ELANe around quasar pairs and signs of filamentary gas connecting two quasars in multiple systems, we conducted the largest study of extended Ly emission around quasar pairs to date. Our sample consists of 14 quasars in eight pairs observed with VLT/MUSE with a snapshot strategy (45 minutes/source). The targets span redshifts of and -band magnitudes between 18 and 22.8, dimmer than most single-quasar samples. Our key findings are as follows:
@@
Twelve out of the 14 observed quasars are associated with extended Ly emission spanning 40 to 160 kpc, and three of the five close pairs are embedded in the same Ly nebula (Table 3, Figs. 3, 7 - 13). In contrast to previously discovered quasar pairs, none of these nebulosities meet the criteria for ELANe in either or extent, lowering the fraction of ELANe around known quasar pairs significantly (Sect. 5.1). On average, the nebulae in our sample are 15 % smaller in maximum projected extent than nebulae around similar single-quasar samples, but they have comparably quiescent kinematics (Section 5.2).
@@
Larger nebulae are preferentially aligned with the expected direction of the large-scale structure between the associated quasar pair, while smaller nebulae tend to be misaligned (Fig. 6). This fact is indicative of a direct connection between filaments and the cool CGM (Sect. 5.3). Wide pairs highly aligned with their filaments, such as quasar pair 4, are excellent targets for ultra-deep follow-up observations to detect these faint IGM structures.
@@
The SB profile of quasar pairs is comparable in brightness to single-quasar profiles at similar redshift but is fainter close to the quasar and falls off more gradually with a power law slope of (Fig. 4). Due to galaxy interactions important in close quasar pairs, the higher surface brightness at larger radii can in part be explained by stripped gas increasing the densities in the CGM. The presence of a second large halo can also add to the Ly emission through increased contribution from IGM structures (Sect. 5.4).
@@
Extended emission around this sample of quasar pairs is morphologically diverse, but on average, it is more asymmetric than single-quasar nebulae with large offsets between the quasar position and the flux-weighted nebula centroid (Table 3, Fig. 5). This can be explained by the presence of more substructures in the overdense regions likely populated by quasar pairs and by close pairs sometimes sharing the same nebula.
@@
We did not find a trend in brightness or morphology of the extended Ly emission with the projected distance between the associated quasar pair, as would be expected if the host galaxies’ interaction influences the CGM in a way that is observable through the extended Ly emission (Sect. 5.4). While the sample studied in this work only encompasses close ( pkpc) and wide ( pkpc) pairs, extending the sample of observed quasar pairs to intermediate projected distances has the potential to reveal the full picture.
@@
This study highlights once again the promise of Ly nebulae around quasars tracing the large-scale structure in dense environments. As quasar pair nebulae show a higher surface brightness level at larger radii than individual quasars, and they provide the most accessible laboratory to directly study the IGM and its connection to the associated galaxies. Indeed, our sample contains excellent targets for future deep and ultra-deep observations to study the large-scale gaseous structures of the IGM (e.g., Tornotti et al. 2024). Additional snapshot observations of quasar pairs are needed to increase the number of such suitable targets and therefore cover the diversity of IGM filament structures and active galaxy interactions at different separation stages.